OpenStudy (anonymous):

The distances (y), in miles, of two cars from their starting points at certain times (x), in hours, are shown by the equations below: Car A: y = 52x + 70 Car B: y = 54x + 56 After how many hours will the two cars be at the same distance from their starting point and what will that distance be? 6 hours, 420 miles 6 hours, 434 miles 7 hours, 420 miles 7 hours, 434 miles

OpenStudy (anonymous):

@Mehek14 @

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

or anyone lol

OpenStudy (anonymous):

@_greatmath7

Mehek (mehek14):

same thing as the other question

Mehek (mehek14):

substitute 52x + 70 for y in the second equation for y

Mehek (mehek14):

\(52x + 70 = 54x + 56\) subtract 54x - 52x = 2x subtract 70 - 56 = 14 \(14=2x\) divide 14/2 = 7 \(7=x\)

Mehek (mehek14):

so we got ts 7 hours not we need to find how many miles

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

C or D

Mehek (mehek14):

\(y=52*7+70\\52*2=364\\364+70=434\)

Mehek (mehek14):

so 7 hrs, 434 mi D

OpenStudy (anonymous):

alright hey im not trying to be annoying but i have more questions.. ugh it kinda sucks tho like i feel like im bothering you

Mehek (mehek14):

nah it's ok

OpenStudy (anonymous):

yay thankss ill post and tag you if i have more to give you more medals